# Equilibrium Selection in Replicator Equations Using Adaptive-Gain Control

### v1

@@ 2.2 Replicator equation
In order to describe the emergent behavior at the population-level, we adopt the standard approach used in population games [1] of considering the population as a continuum of players of unit mass. Each player repeatedly engages in the two-player matrix game described above with their peers, revising their strategy toward improving their payoff. Specifically, we denote by $x(t)\in[0,1]$ the fraction of adopters of action $1$ at time $t\geq 0$; consequently, the fraction of adopters of action $2$ is equal to $1-x(t)$. Then, the total reward $r_{i}(x,\boldsymbol{A})$ associated with action $i$ is given by the average payoff that a player receives for choosing that action from all games played [1], which is equal to
@@ 3.1 Adaptive-gain controller
$\begin{array}[]{l}\boldsymbol{G^{(1)}}:=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},\,\,\boldsymbol{G^{(2)}}:=\begin{bmatrix}0&1\\ 0&0\end{bmatrix},\,\,\boldsymbol{G^{(3)}}:=\begin{bmatrix}0&0\\ 1&0\end{bmatrix},\,\,\boldsymbol{G^{(4)}}:=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}.\end{array}$ (6)
@@
The regularity of $\phi$ and the presence of the term $g$ on the right-hand side of the second equation of Eq. (8) guarantees that $g(t)\in\mathbb{R}_{+}$ for all $t\in\mathbb{R}_{+}$. This observation, combined with the regularity of the right-hand side of the first equation of Eq. (8) and the fact that it is equal to 00 at the boundaries $x=0$ and $x=1$ yields the claim [52]. ∎
@@ 3.2 Problem formulation
In case i), consider $\bar{x}=0$ when $b-d>0$. Since $b-d>0$, then, by continuity, there exist two positive constants $\delta,\gamma>0$ such that $(a+d-b-c)x+b-d>\gamma$, for all $x\in[0,\delta]$. The hypothesis that $\lim_{t\to\infty}g(t)=0$ implies that there exists $\bar{t}$ such that $g(t)\leq{\gamma}$ for all $t\geq\bar{t}$. As a consequence, we can bound $\dot{x}(t)\geq\frac{1}{2}\gamma x(1-x)>0$ for any $t\geq\bar{t}$ and $x\in[0,\delta]$. Note that, regularity of $\phi(x)$ implies that exists a constant $\bar{\gamma}$ such that $g(t)\geq\bar{\gamma}$ for any $t\in[0,\bar{t}]$. Therefore, from Eq. (8) we bound $\dot{x}\geq-(\max\{G_{12},G_{22}\}\bar{\gamma}-c-d)x$ for $t\in[0,\bar{t}$), which implies, due to Gronwall’s inequality [40], that $x(\bar{t})\geq x(0)\exp\{-(\max\{G_{12},G_{22}\}\bar{\gamma}-c-d)\bar{t}\}$ in the same interval. In other words, $x(t)$ cannot converge to $\bar{x}=0$ before time $\bar{t}$ and, for any $t\geq\bar{t}$, $\dot{x}(t)$ is strictly positive in a neighborhood $[0,\delta]$ of $\bar{x}=0$. This latter property implies that $x(t)$ cannot converge to 00 after $\bar{t}$. A similar argument can be used for $\bar{x}=1$ when $c-a>0$.
@@ 4.1 Conformity gain controller
$\begin{bmatrix}-3x^{2}(\alpha+\beta+g)+2x(\alpha+2\beta+g)-\beta-g&-x(1-x)^{2}\\ g\phi^{\prime}(x)&\phi(x)\end{bmatrix}.$
In fact, at $(0,0)$, we get two negative eigenvalues $-\beta$ and $\phi(0)<0$; at $(1,0)$ we get the eigenvalues $-\alpha-\beta$ and $\phi(1)>0$; at $(\frac{\beta}{\alpha+\beta},0)$, we get the eigenvalues $\frac{\alpha\beta}{\alpha+\beta}>0$ and $\phi(\frac{\beta}{\alpha+\beta})>0$, where the latter holds true because $\frac{\beta}{\alpha+\beta}$ is the unstable equilibrium of the uncontrolled dynamics. Assumption 1 implies that $\frac{\beta}{\alpha+\beta}>\delta$ and, due to condition 2 of the theorem, $\phi(\frac{\beta}{\alpha+\beta})>0$.
@@
$\dot{x}=x(1-x)((\alpha+\beta)x-\beta-g(1-x))\leq x(1-x)(\alpha-\varepsilon g(0)e^{\mu t}),$ (11)
@@
In the third step, we rule out convergence to $\mathcal{C}$. Let us assume that there exists a trajectory of Eq. (9) with initial condition in $\mathcal{B}$ that converges to $x=1$. Hence, for any constant $\varepsilon>0$ there exists $\bar{t}$ such that $x(t)>1-\varepsilon$, for any $t>\bar{t}_{\varepsilon}$. Let us consider $\varepsilon=\varepsilon^{*}$ from the third condition of the theorem on $\phi(x)$. This implies that, for any $t>\bar{t}_{\varepsilon}$ it holds $\phi(x)>\alpha$. Moreover, since $\phi(x)>\alpha$ for all $x$ in the compact domain $[1-\varepsilon^{*},1]$, we can refine the bound as $\phi(x)\geq\alpha+\sigma$, for some constant $\sigma>0$. Hence, similar to the above, we can bound the second equation of Eq. (9) as $\dot{g}\geq(\alpha+\sigma)g$ for any $t\geq\bar{t}_{\varepsilon}$. As a consequence, for $t\geq\bar{t}_{\varepsilon}$, there holds
@@
being $\dot{g}>0$ in $\mathcal{B}\cup\mathcal{C}$. On the other hand, if we define $y=1-x$, and notice that $\dot{y}=-\dot{x}$, it follows from Eq. (9) that
@@
Condition 3) in Theorem 1 requires some information on the order of magnitude of the payoff function. However, it is worth noticing that one does not need full knowledge of the payoff matrix in order to guarantee that Condition 3) is satisfied. In fact, one only needs an estimate of the difference between the entries $a-c$ or, conservatively, just an upper bound on the maximum component of the payoff matrix, which can be ultimately estimated online, from observations of the velocity of the dynamics. For instance, when using Eq. (17), it suffices to guarantee that $p$ is larger than (an estimate of) the maximum entry of $\boldsymbol{A}$.
@@ 4.2 Innovation gain controller
$\dot{x}(t)$ $\leq x(t)(1-x(t))\left((\alpha+\beta)x(t)-\beta-x(t)g(0)e^{\mu t}\right)$ (19)
@@ 4.3 Comparison between the two control approaches
From the simulations in Fig. 3, we observe that both our controllers successfully steer the system to the desired equilibrium $\bar{x}=0$. These simulations give us also the opportunity to discuss some of the pros and cons of the two approaches. Specifically, we observe that the conformity gain control appears to yield faster convergence. This is due to the fact that the gain is multiplied by $1-x(t)$ instead of by $x(t)$ in the $\dot{x}$ equation. Therefore, its effect does not vanish as $x(t)$ approaches 00. The drawback, however, is that larger peaks in the gain are reached when the system is far from the desired equilibrium, as can be observed in Fig. 3a.
@@
[…]
@@ 5 Consensus Stabilization Problem
whose right-hand side is the bound from Eq. (21). For Eq. (22), $\bar{y}=0$ has the same linear stability properties of $\bar{x}=0$ for the uncontrolled system in Eq. (4), determined by linearization. Therefore, it is unstable and it cannot be attractive. Finally, Gronwall’s inequality on Eq. (21), we conclude that $x(t)\geq y(t)$, solution of Eq. (22) with initial condition $y(0)=x(0)$. Since $y(t)$ does not converge to 00, necessarily $x(t)$ cannot converge to 00 too, yielding the claim. ∎
@@
To obtain a contradiction, assume that $g(t)\to\infty$. First, we exclude the possibility that $x(t)$ converges to $1$. This is done by following the same line of reasoning used in the proof of Theorem 1. In fact, due to condition 3) of the Theorem’s statement, $g(t)$ grows in a neighborhood of $x=1$ faster than an exponentially-growing function with exponent $\alpha$. Hence, $g(t)>\beta$ in finite time $\tau_{1}$, yielding $\dot{x}\leq\alpha x(1-x)$ for any $t\geq\tau_{1}$. Consequently, $x(t)\to 1$ at an exponential rate no larger than $\alpha$. Combining this with the fact that $g(t)$ grows exponentially at a rate larger than $\alpha$, it follows that the term $g(1-x)$ grows exponentially large, guaranteeing that there exists a time $\tau_{2}$ such that $\dot{x}(t)<0$ for all $t\geq\tau_{2}$, and thus $x(t)$ eventually decreases and cannot converge to $1$.
Since $x(t)$ does not converge to $1$ and we have assumed that $g(t)\to\infty$, then there exist two positive constants $\varepsilon_{1}>0$ and $\tau_{3}<\infty$ such that i) $x(\tau_{3})\leq 1-\varepsilon_{1}$ and ii) $g(\tau_{3})\geq\alpha/\varepsilon_{1}+\beta+\gamma$, where $\gamma>0$ is a sufficiently small but positive constant. This implies that $\dot{x}(\tau_{3})\leq-\gamma x(1-x)^{2}$. Using again the continuity argument used in the proof of Theorem 1 and the fact that $g(t)$ is monotonically increasing, we conclude that the bound $\dot{x}(t)\leq-\gamma\varepsilon_{1}^{2}x$ holds true for any $t\geq\tau_{3}$. Hence, we use Gronwall’s lemma to derive an exponential bound on the convergence of $x(t)$ to 00 as
@@
[…]
@@
[…]
@@
The analysis is similar to the one of the dominant-strategy game in Eq. (23). Specifically, we proceed by contradiction, showing that the limit of $g(t)$ exists finite and thus $x(t)$ must converge to 00. It worth noticing that the negative sign in the first equation of Eq. (27) simplifies the treatment. In fact, $\dot{x}$ is always negative in a neighborhood of $x=1$, regardless of $g(t)$. This implies that convergence to $x=1$ is excluded without needing any additional condition on $\phi$. The rest of the proof then follows the identical line of arguments used for the dominant-strategy game, and detailed calculations are omitted. ∎
@@ 6.1 Dominant-strategy games
$\dot{x}$ $=x(1-x)(\alpha x+\beta(1-x)-\bar{g}x)-x^{2}(1-x)(g-\bar{g})$ (32)
@@
Consequently, it holds true that $V(x,g)>0$ for all $x\in[0,1]$ and $g\geq 0$, and $V(x,g)=0$ if and only if $x=\bar{x}$ and $g=\bar{g}$. Finally, we observe that $V$ diverges to +$\infty$+ ∞ at the boundaries of the domain $[0,1]\times[0,\infty]$. These properties make $V(x,g)$ a suitable Lyapunov candidate function.

### v2

@@ 2.2 Replicator equation
In order to describe the emergent behavior at the population-level, we adopt the standard approach used in population games [1] of considering the population as a continuum of players of unit mass. Each player repeatedly engages in the two-player matrix game described above with their peers, revising their strategy toward improving their payoff. Specifically, we denote by $x(t)\in[0,1]$ the fraction of adopters of action $1$ at time $t\geq 0$; consequently, the fraction of adopters of action $2$ is equal to 1x(t)1-x(t). Then, the total reward $r_{i}(x,\boldsymbol{A})$ associated with action $i$ is given by the average payoff that a player receives for choosing that action from all games played [1], which is equal to
@@ 3.1 Adaptive-gain controller
[…]
@@
The regularity of $\phi$ and the presence of the term $g$ on the right-hand side of the second equation of Eq. (8) guarantees that $g(t)\in\mathbb{R}_{+}$ for all $t\in\mathbb{R}_{+}$. This observation, combined with the regularity of the right-hand side of the first equation of Eq. (8) and the fact that it is equal to 0 at the boundaries $x=0$ and $x=1$ yields the claim [52]. ∎
@@ 3.2 Problem formulation
In case i), consider $\bar{x}=0$ when bd>0b-d>0. Since bd>0b-d>0, then, by continuity, there exist two positive constants $\delta,\gamma>0$ such that $(a+d-b-c)x+b-d>\gamma$, for all $x\in[0,\delta]$. The hypothesis that $\lim_{t\to\infty}g(t)=0$ implies that there exists $\bar{t}$ such that $g(t)\leq{\gamma}$ for all $t\geq\bar{t}$. As a consequence, we can bound $\dot{x}(t)\geq\frac{1}{2}\gamma x(1-x)>0$ for any $t\geq\bar{t}$ and $x\in[0,\delta]$. Note that, regularity of $\phi(x)$ implies that exists a constant $\bar{\gamma}$ such that $g(t)\geq\bar{\gamma}$ for any $t\in[0,\bar{t}]$. Therefore, from Eq. (8) we bound $\dot{x}\geq-(\max\{G_{12},G_{22}\}\bar{\gamma}-c-d)x$ for $t\in[0,\bar{t}$), which implies, due to Gronwall’s inequality [40], that $x(\bar{t})\geq x(0)\exp\{-(\max\{G_{12},G_{22}\}\bar{\gamma}-c-d)\bar{t}\}$ in the same interval. In other words, $x(t)$ cannot converge to $\bar{x}=0$ before time $\bar{t}$ and, for any $t\geq\bar{t}$, $\dot{x}(t)$ is strictly positive in a neighborhood $0,\delta]$ of $\bar{x}=0$. This latter property implies that $x(t)$ cannot converge to 0 after $\bar{t}$. A similar argument can be used for $\bar{x}=1$ when ca>0c-a>0.
@@ 4.1 Conformity gain controller
$\begin{bmatrix}-3x^{2}(\alpha+\beta+g)\hskip-1.42271pt+\hskip-1.42271pt2x(\alpha+2\beta+g)\hskip-1.42271pt-\hskip-1.42271pt\beta\hskip-1.42271pt-\hskip-1.42271ptg&-x(1-x)^{2}\\ g\phi^{\prime}(x)&\phi(x)\end{bmatrix}\hskip-2.84544pt$.
In fact, at $(0,0)$, we get two negative eigenvalues $\beta$ and $\phi(0)<0$; at $(1,0)$ we get the eigenvalues $\alpha-\beta$ and $\phi(1)>0$; at $(\frac{\beta}{\alpha+\beta},0)$, we get the eigenvalues $\frac{\alpha\beta}{\alpha+\beta}>0$ and $\phi(\frac{\beta}{\alpha+\beta})>0$, where the latter holds true because $\frac{\beta}{\alpha+\beta}$ is the unstable equilibrium of the uncontrolled dynamics. Assumption 1 implies that $\frac{\beta}{\alpha+\beta}>\delta$ and, due to condition 2 of the theorem, $\phi(\frac{\beta}{\alpha+\beta})>0$.
@@
$\dot{x}=x(1-x)((\alpha+\beta)x-\beta-g(1-x))\hskip-1.42271pt\leq\hskip-1.42271ptx(1-x)(\alpha-\varepsilon g(0)e^{\mu t}),$ (11)
@@
In the third step, we rule out convergence to $\mathcal{C}$. Let us assume that there exists a trajectory of Eq. (9) with initial condition in $\mathcal{B}$ that converges to $x=1$. Hence, for any constant $\varepsilon>0$ there exists $\bar{t}$ such that $x(t)>1-\varepsilon$, for any $t>\bar{t}_{\varepsilon}$. Let us consider $\varepsilon=\varepsilon^{*}$ from the third condition of the theorem on $\phi(x)$. This implies that, for any $t>\bar{t}_{\varepsilon}$ it holds $\phi(x)>\alpha$. Moreover, since $\phi(x)>\alpha$ for all $x$ in the compact domain $1-\varepsilon^{*},1]$, we can refine the bound as $\phi(x)\geq\alpha+\sigma$, for some constant $\sigma>0$. Hence, similar to the above, we can bound the second equation of Eq. (9) as $\dot{g}\geq(\alpha+\sigma)g$ for any $t\geq\bar{t}_{\varepsilon}$. As a consequence, for $t\geq\bar{t}_{\varepsilon}$, there holds
@@
being $\dot{g}>0$ in $\mathcal{B}\cup\mathcal{C}$. On the other hand, if we define y=1xy=1-x, and notice that $\dot{y}=-\dot{x}$, it follows from Eq. (9) that
@@
Condition 3) in Theorem 1 requires some information on the order of magnitude of the payoff function. However, it is worth noticing that one does not need full knowledge of the payoff matrix in order to guarantee that Condition 3) is satisfied. In fact, one only needs an estimate of the difference between the entries aca-c or, conservatively, just an upper bound on the maximum component of the payoff matrix, which can be ultimately estimated online, from observations of the velocity of the dynamics. For instance, when using Eq. (17), it suffices to guarantee that $p$ is larger than (an estimate of) the maximum entry of $\boldsymbol{A}$.
@@ 4.2 Innovation gain controller
$\dot{x}(t)$ $\leq x(t)(1-x(t))\left((\alpha+\beta)x(t)\hskip-1.42271pt-\hskip-1.42271pt\beta\hskip-1.42271pt-\hskip-1.42271ptx(t)g(0)e^{\mu t}\right)$ (19)
@@ 4.3 Comparison between the two control approaches
From the simulations in Fig. 3, we observe that both our controllers successfully steer the system to the desired equilibrium $\bar{x}=0$. These simulations give us also the opportunity to discuss some of the pros and cons of the two approaches. Specifically, we observe that the conformity gain control appears to yield faster convergence. This is due to the fact that the gain is multiplied by 1x(t)1-x(t) instead of by $x(t)$ in the $\dot{x}$ equation. Therefore, its effect does not vanish as $x(t)$ approaches 0. The drawback, however, is that larger peaks in the gain are reached when the system is far from the desired equilibrium, as can be observed in Fig. 3a.
@@
Besides integrating the two different approaches, the constraints in our theoretical results leave many other degrees of freedom in designing the adaptation rate $\phi$. Hence, the optimal design of such a function is of paramount importance toward improving the performance of our controller. To investigate this issue, we perform a set of simulations, in which we numerically integrate Eq. (8) over a time horizon $T$, using the innovation gain control and different adaptation rates in the form of Eq. (17), exploring how the parameters $p$ and $q$ affect the total control effort $J_{g}:=\int_{0}^{T}g(t)dt$ and the maximum gain $g_{\text{max}}:=\max_{t\in[0,T]}g(t)$. Our results, reported in Fig. 4, depicts an interesting scenario. In fact, while the total control effort displays a monotonic behavior, whereby it grows as $q$ increase and $p$ decreases, the peak gain shows a more complex, non-monotonic behavior. Specifically, $g_{\text{max}}$ seems to be particularly large when $p$ is large and $q$ has moderate values, while its minimum is achieved in an intermediate region for the parameter $p$, which depends on $q$. Apparently, an optimal trade-off can be obtained by setting moderate levels of $p$ and small values of $q$, similar to Eq. (3b). Alternatively, for larger values of $q$, the plot in Fig. 4b seems to describe a linear relation between the parameters $q$ and $p$ that minimizes the maximum peak gain, whereby the optimal value of $p$ (initially small) increases linearly with $q$.
@@ 5 Consensus Stabilization Problem
whose right-hand side is the bound from Eq. (21). For Eq. (22), $\bar{y}=0$ has the same linear stability properties of $\bar{x}=0$ for the uncontrolled system in Eq. (4), determined by linearization. Therefore, it is unstable and it cannot be attractive. Finally, Gronwall’s inequality on Eq. (21), we conclude that $x(t)\geq y(t)$, solution of Eq. (22) with initial condition $y(0)=x(0)$. Since $y(t)$ does not converge to 0, necessarily $x(t)$ cannot converge to 0 too, yielding the claim. ∎
@@
To obtain a contradiction, assume that $g(t)\to\infty$. First, we exclude the possibility that $x(t)$ converges to $1$. This is done by following the same line of reasoning used in the proof of Theorem 1. In fact, due to condition 3) of the Theorem’s statement, $g(t)$ grows in a neighborhood of $x=1$ faster than an exponentially-growing function with exponent $\alpha$. Hence, $g(t)>\beta$ in finite time $\tau_{1}$, yielding $\dot{x}\leq\alpha x(1-x)$ for any $t\geq\tau_{1}$. Consequently, $x(t)\to 1$ at an exponential rate no larger than $\alpha$. Combining this with the fact that $g(t)$ grows exponentially at a rate larger than $\alpha$, it follows that the term g(1x)g(1-x) grows exponentially large, guaranteeing that there exists a time $\tau_{2}$ such that $\dot{x}(t)<0$ for all $t\geq\tau_{2}$, and thus $x(t)$ eventually decreases and cannot converge to $1$.
Since $x(t)$ does not converge to $1$ and we have assumed that $g(t)\to\infty$, then there exist two positive constants $\varepsilon_{1}>0$ and $\tau_{3}<\infty$ such that i) $x(\tau_{3})\leq 1-\varepsilon_{1}$ and ii) $g(\tau_{3})\geq\alpha/\varepsilon_{1}+\beta+\gamma$, where $\gamma>0$ is a sufficiently small but positive constant. This implies that $\dot{x}(\tau_{3})\leq-\gamma x(1-x)^{2}$. Using again the continuity argument used in the proof of Theorem 1 and the fact that $g(t)$ is monotonically increasing, we conclude that the bound $\dot{x}(t)\leq-\gamma\varepsilon_{1}^{2}x$ holds true for any $t\geq\tau_{3}$. Hence, we use Gronwall’s lemma to derive an exponential bound on the convergence of $x(t)$ to 0 as
@@
Observe that $M:=\max_{x\in[0,1]}\phi(x)<\infty$, being the maximum of a continuous function over a compact set. Hence, we are ready to obtain the last piece of our proof by contradiction, by integrating the second equation in Eq. (23), obtaining $g(t)=g(0)e^{k\int_{0}^{t}\phi(x(s))ds}$. Finally, by using that $\phi(x(t))\leq M$ for all $t\leq\tau_{4}$ and
@@
$\lim_{t\to\infty}g(t)$ $=\lim_{t\to\infty}g(0)e^{\int_{0}^{t}\phi(x(s))ds}$ (26)
$=g(0)e^{\int_{0}^{\tau_{4}}\phi(x(s))ds}e^{\int_{\tau_{4}}^{\infty}\phi(x(s))ds}$
$\leq g(0)e^{\int_{0}^{\tau_{4}}Mds}e^{k(1-\varepsilon_{1})^{h}\int_{\tau_{4}}^{\infty}e^{-h\gamma\varepsilon_{1}^{2}(t-\tau_{4})}ds}$
$\leq g(0)e^{M\tau_{4}}e^{k(1-\varepsilon)^{h}\int_{0}^{\infty}e^{-h\gamma\varepsilon_{1}^{2}s}ds}$
@@
The analysis is similar to the one of the dominant-strategy game in Eq. (23). Specifically, we proceed by contradiction, showing that the limit of $g(t)$ exists finite and thus $x(t)$ must converge to 0. It worth noticing that the negative sign in the first equation of Eq. (27) simplifies the treatment. In fact, $\dot{x}$ is always negative in a neighborhood of $x=1$, regardless of $g(t)$. This implies that convergence to $x=1$ is excluded without needing any additional condition on $\phi$. The rest of the proof then follows the identical line of arguments used for the dominant-strategy game, and detailed calculations are omitted. ∎
@@ 6.1 Dominant-strategy games
$\dot{x}$ $=x(1-x)(\alpha x+\beta(1-x)\hskip-1.42271pt-\hskip-1.42271pt\bar{g}x)\hskip-1.42271pt-\hskip-1.42271ptx^{2}(1\hskip-1.42271pt-\hskip-1.42271ptx)(g-\bar{g})$ (32)
@@
Consequently, it holds true that $V(x,g)>0$ for all $x\in[0,1]$ and $g\geq 0$, and $V(x,g)=0$ if and only if $x=\bar{x}$ and $g=\bar{g}$. Finally, we observe that $V$ diverges to $\infty$ at the boundaries of the domain $0,1]\times[0,\infty]$. These properties make $V(x,g)$ a suitable Lyapunov candidate function.